Diffusion morning report 02/27/06

Ted Berg tberg at snf.stanford.edu
Mon Feb 27 07:14:16 PST 2006


Good Morning All,
    Looking like Monday this morning. Epi is down for Temp issue. Poly 
has pressure control issues. LTO has possible temp control
issues. LTO is already way over 30microns-someone will be bumped so we 
can do pull and clean. Everything else is OK for now.


>   	Date: 	2/27/2006 	THICKNESS-Change at
> Tube Name/Number 	Status 	Comments 	 
> Tylan 1 	Up 	  	 
> Tylan 2 	Up 	  	 
> Tylan 3 	Up 	  	 
> Tylan 4 	Up 	  	 
> Tylan 5-BBr3 	Up 	  	 
> Tylan 6-POCl3 	Up 	  	 
> Tylan 7-TEOS2 	Up 	  	 
> Tylan 8-SiGe 	Up 	  	 18.99/30
> Tylan 9-Poly 	Up 	  	3.44/30
> Tylan 10-Nitride 	Up 	  	8.24/30
> Tylan 11-BPSG-LTO 	Up 	  	37.9/30
> Tylan 12-FGA 	Up 	  	 
>   	  	  	 
> Thermco 1 	Up 	  	 
> Thermco 2 	Up 	  	 
> Thermco Poly 	Up 	  	 
> Thermco Nitride 	Down 	Testing 	 
>   	  	  	 
> Burn Box 	Up 	  	 
> ZYGO 	up 	  	  
>   	  	  	 
> GAAS 22 	UP 	  	 
> GAAS 23 	up 	  	 
>   	  	  	 
> SiH4 for furnaces 	  	A on line 120 	 changeover working well 
>

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060227/2cfa100e/attachment.html>


More information about the diffusion mailing list